# Cross-correlation between the curvature perturbations and magnetic fields in pure ultra slow roll inflation

### v1

@@ I Introduction
Large-scale, coherent magnetic fields are observed in galaxies, clusters of galaxies, and even in the intergalactic medium (for reviews on magnetic fields, see Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 10]). The Fermi/LAT and HESS observations of TeV blazars suggest that the strength of the magnetic fields in the intergalactic voids is of the order of $10^{-16}\,\mathrm{G}$ [11, 12, 13, 14, 15, 16, 17]. It may not be possible to explain the presence of magnetic fields in the intergalactic voids solely on the basis of astrophysical phenomena such as batteries [3, 4]. Hence, it seems necessary to invoke a cosmological mechanism for the primordial origin of these fields. While the Fermi/LAT and HESS observations provide a lower bound on the current strengths of the cosmological magnetic fields, the anisotropies in the cosmic microwave background (CMB) provide an upper bound. The primordial magnetic fields can induce scalar, vector, and tensor perturbations, which, in turn, can leave distinct signatures on the temperature and polarization angular power spectra of the CMB [18, 19, 20, 21, 22, 23, 24]. The CMB observations lead to an upper bound on the primordial magnetic fields to be of the order of $10^{-9}\,\mathrm{G}$ [25, 26, 27, 28, 29, 30, 31].
@@
This manuscript is organized as follows. In the following section, we shall discuss the behavior of the background and the Fourier modes describing the curvature perturbations in specific pure ultra slow roll scenarios. We shall also outline the generation of electromagnetic fields by a non-conformal coupling function that depends on the inflaton and discuss the behavior of the Fourier modes describing the electromagnetic vector potential. Further, we shall arrive at the power spectra of the curvature perturbations and the electromagnetic fields in these cases. In Sec. III, we shall consider a non-conformal coupling function that depends on the kinetic energy of the inflaton. We shall show that, for suitable choices of the parameters involved, such a coupling function leads to the desired scale invariant spectrum for the magnetic field in the ultra slow roll scenarios. In Sec. IV, we shall derive the third order action that describes the interaction between the curvature perturbation and the electromagnetic field for the new choice of the non-conformal coupling function. We shall also arrive at formal expressions for the contributions to the three-point cross-correlation (in Fourier space) between the curvature perturbations and the magnetic fields corresponding to the different terms in the action at the cubic order. We shall also introduce the non-Gaussianity parameter associated with the three-point cross-correlation of interest. In Sec. V, we shall calculate the three-point cross-correlation for two specific cases of pure ultra slow roll inflation and suitable values of the relevant parameters. We shall also illustrate the complete structure of the non-Gaussianity parameter and discuss the validity of the consistency condition governing the parameter in the scenarios we consider. Lastly, in Sec. VI, we shall conclude with a summary and outlook. We relegate complementary details of the computation and arguments on certain related issues to the appendices.
@@ II Background, mode functions and power spectra
In this section, we shall discuss the behavior of the background as well as the Fourier modes of the curvature perturbation and the electromagnetic field in ultra slow roll inflation. We shall also discuss the power spectra of the curvature perturbations and the electromagnetic fields that arise in a couple of specific situations.
@@ II.1 Evolution of the background
where $H=\dot{a}/a$ is the Hubble parameter and $V_{\phi}={\mathrm{d}}V/{\mathrm{d}}\phi$. As we mentioned above, a brief epoch of ultra slow roll inflation is usually realized in potentials that contain a point of inflection, i.e. a point wherein the first and the second derivatives of the potential, viz. $V_{\phi}$ and $V_{\phi\phi}$, vanish (see, for instance, Refs. [60, 61, 48]). Note that, when $V_{\phi}=0$, the equation of motion for the scalar field reduces to
@@
where $\epsilon_{1\mathrm{i}}$ is the initial value of the first slow roll parameter when the value of the scale factor is $a_{\mathrm{i}}=-1/(H_{{}_{\mathrm{I}}}\,\eta_{\mathrm{i}})$ at the conformal time $\eta_{\mathrm{i}}$. In our discussion below, we shall consider two cases wherein $p=6$ and $p=4$. As we shall see, the first of these leads to a scale invariant spectrum for the curvature perturbations, while the second leads to simple mode functions, which allow us to easily calculate the two and three-point functions of interest.
@@ II.2 Mode functions describing the curvature perturbation and the scalar power spectrum
There are two points concerning this scalar power spectrum that need to be emphasized. The first point is the fact the spectrum is independent of $k$, i.e. it is scale invariant. The second point is that, in contrast to slow roll inflation, the spectrum does not approach a constant value at late times [i.e. when $k\ll\sqrt{z^{\prime\prime}/z}$, which is equivalent to the super-Hubble limit $k/(a\,H)\simeq(-k\,\eta)\ll 1$]. In fact, the power spectrum grows as $a^{6}$ at late times. This can be attributed to the fact that the mode function $f_{k}$ grows as $a^{3}$ in the super-Hubble limit. Such a growth is a well known feature in ultra slow roll inflationary scenarios [55, 56].
@@ II.3 Mode functions describing the electromagnetic field and power spectra
We can arrive at the equation of motion governing the vector potential $A_{\mu}$ by varying the above action. In what follows, we shall choose to work in the Coulomb gauge wherein $A_{0}=0$ and $(g^{ij}\,\partial_{i}A_{j})=0$. Let $\bar{A}_{k}$ denote the Fourier modes associated with the vector potential $A_{i}$. We find that the mode functions $\bar{A}_{k}$ satisfy the following differential equation (see, for example, Refs. [35, 5]):
@@
where $|0\rangle$ denotes the vacuum state annihilated by the annihilation operator $\hat{b}_{{\bm{k}}}^{\lambda}$, i.e. $\hat{b}_{{\bm{k}}}^{\lambda}|0\rangle=0$ for all ${\bm{k}}$ and $\lambda$. On utilizing the decomposition (20) of the vector potential, the power spectra $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ and $\mathcal{P}_{{}_{\mathrm{E}}}(k)$ can be expressed in terms of the mode functions $\bar{A}_{k}$ and their scaled forms ${\mathcal{A}}_{k}$ as follows [35, 5]:
@@
where $\eta_{\mathrm{e}}$ is the conformal time at the end of inflation and $\bar{n}$ is a number. Note that the overall coefficient has been chosen so that the coupling function $J$ reduces to unity at the end of inflation. In situations wherein the scale factor can be expressed in the de Sitter form, the coupling function is given by
@@
For such a choice of the coupling function, the solution to Eq. (18) that satisfies the Bunch-Davies initial conditions in the domain wherein $k\gg\sqrt{J^{\prime\prime}/J}$ (which corresponds to the sub-Hubble limit $(-k\,\eta)\ll 1$ for the $J$ we are working with) can be expressed as
@@
The spectra of the electromagnetic fields can be evaluated at late times when $k\ll\sqrt{J^{\prime\prime}/J}$, i.e. in super-Hubble limit wherein $(-k\,\eta)\ll 1$. Upon substituting the mode functions (27) in the expressions (24) and considering the super-Hubble limit, the spectra of the magnetic and electric fields $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ and $\mathcal{P}_{{}_{\mathrm{E}}}(k)$ can be obtained to be [35, 5]
@@
$m=\begin{cases}\bar{n},&\text{for $\bar{n}<-\frac{1}{2}$},\\ -\bar{n}-1,&\text{for $\bar{n}>-\frac{1}{2}$}.\end{cases}$ (30)
@@
$m=\begin{cases}\bar{n},&\text{for $\bar{n}<\frac{1}{2}$},\\ 1-\bar{n},&\text{for $\bar{n}>\frac{1}{2}$}.\end{cases}$ (31)
@@
There are a few points that we need to clarify regarding the results that we have arrived at above. To begin with, note that, we obtain a scale invariant spectrum for the magnetic field when $\bar{n}=2$ and when $\bar{n}=-3$. However, when $\bar{n}=-3$, the spectrum of the electric field behaves as $\mathcal{P}_{{}_{\mathrm{E}}}(k)\propto(-k\,\eta_{\mathrm{e}})^{-2}$, which grows indefinitely at late times. In other words, the energy in the electric field becomes large leading to significant backreaction on the background (in this regard, see, for instance, Refs. [38, 53]). Therefore, when we require a scale invariant spectrum for the magnetic field, we shall focus on the $\bar{n}=2$ case. Secondly, recall that the original non-conformal coupling function $J$ depends on $\phi$ [cf. Eq. (16)]. Whereas, our choices for $J$ in Eqs. (25) and (26) depend on the scale factor and conformal time. Such a behavior for $J(\eta)$ is achievable in most models of inflation that permit slow roll [37, 36, 74]. However, there exist situations, in particular, when an inflationary potential permits a period of ultra slow roll, wherein a behavior such as $J\propto a^{2}\propto\eta^{-2}$ may not be achievable (for discussions in this context, see Ref. [53]). To circumvent such challenges, either, one has to consider inflation involving more than one field (for a discussion on this point, see Ref. [54]) or, as we shall discuss in the following section, consider more non-trivial forms for the coupling functions involving the kinetic energy of the inflaton.
@@ III Construction of the non-conformal coupling function in pure ultra slow roll inflation
where $J_{0}$ is a constant which we shall choose so that $J$ reduces to unity at the end of inflation. It should be evident that, since the non-conformal coupling function $J(X)$ depends only on time, the electromagnetic field described by the action (34) can be treated in the same manner as in the case of the field described by the earlier action (16). For instance, if we work in the Coulomb gauge, the Fourier modes $\bar{A}_{k}$ of the vector potential $A_{i}$ will continue to satisfy the same equation of motion, viz. Eq. (17). Also, on quantization, we can elevate the vector potential $A_{i}$ to the operator $\hat{A}_{i}$ as in Eq. (20). Note that, in the ultra slow roll inflationary scenarios of interest, we can write the coupling function (35) as
@@
On varying the action (34) with respect to the metric tensor, we obtain the stress-energy tensor for the electromagnetic field to be
@@
We find that, in the background described by the FLRW line-element (1), the energy density of the electromagnetic field, viz. $\rho_{{}_{\mathrm{EM}}}=-T^{t}_{t}$, can be expressed as
@@
Using the above solutions in Eqs. (24), we obtain the power spectra of the magnetic and electric fields to be
@@ IV.1 Action at the cubic order
As is often done, we shall compute the three-point function using the standard methods of perturbative quantum field theory (for calculations involving scalars and tensors, see, for example, Refs. [94, 95, 96, 97, 98], and for discussions involving magnetic fields, see Refs. [87, 88, 89, 90, 91, 92, 99, 93]). Our first task is to obtain the interaction Hamiltonian at the cubic order that governs the cross-correlation of interest. We shall start with the action (34), take into account the scalar perturbations in the spatially flat, FLRW line-element, and arrive at the cubic order action involving the curvature perturbations and the electromagnetic fields. We shall make use of the Arnowitt–Deser–Misner (ADM) formalism to obtain the action [100]. Recall that, in the ADM formalism, the line-element associated with a generic spacetime is described in terms of the lapse function $N$, the shift vector $N^{i}$ and the spatial metric $g_{ij}$ as follows:
@@
Note that the metric elements $g_{00}$ and $g_{0i}$ associated with the above line-element are given by
@@
$N_{1}=\frac{{\mathcal{R}}^{\prime}}{a\,H},\;\chi=-\frac{{\mathcal{R}}}{a\,H}+\frac{\epsilon_{1}}{a}\,\nabla^{-2}{\mathcal{R}}^{\prime}.$ (45)
On substituting the above components of the metric tensor in the action (34), we find that the action up to the cubic order involving the curvature perturbation ${\mathcal{R}}$ and the electromagnetic vector potential $A_{i}$ can be expressed as
@@
There are two clarifying remarks that we wish to make regarding the interaction Hamiltonian $H_{\mathrm{int}}$ we have obtained above. Firstly, note that, in the above Hamiltonian, if we ignore the terms containing $n$ within the curly brackets, it reduces to the standard form associated with the original action (16) involving $J(\phi)$ (in this regard, see, for example, Refs. [88, 89]). Secondly, as with any action at the cubic order, the interaction Hamiltonian proves to be negative of the corresponding cubic order Lagrangian.

### v2

@@ I Introduction
Large-scale, coherent magnetic fields are observed in galaxies, clusters of galaxies, and even in the intergalactic medium (for reviews on magnetic fields, see Refs. [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 10]). The Fermi/LAT and HESS observations of TeV blazars suggest that the strength of the magnetic fields in the intergalactic voids is of the order of $10^{-16}\,\mathrm{G}$ [11, 12, 13, 14, 15, 16, 17]. It may not be possible to explain the presence of magnetic fields in the intergalactic voids solely on the basis of astrophysical phenomena such as batteries (see Refs. [3, 4]; however, in this context, see Refs. [18, 19]). Hence, it seems necessary to invoke a cosmological mechanism for the primordial origin of these fields. While the Fermi/LAT and HESS observations provide a lower bound on the current strengths of the cosmological magnetic fields, the anisotropies in the cosmic microwave background (CMB) provide an upper bound. The primordial magnetic fields can induce scalar, vector, and tensor perturbations, which, in turn, can leave distinct signatures on the temperature and polarization angular power spectra of the CMB [20, 21, 22, 23, 24, 25, 26]. The CMB observations lead to an upper bound on the primordial magnetic fields to be of the order of $10^{-9}\,\mathrm{G}$ [27, 28, 29, 30, 31, 32, 33].
@@
This manuscript is organized as follows. In the following section, we shall discuss the behavior of the background and the Fourier mode functions describing the curvature perturbations in specific pure ultra slow roll scenarios. We shall also outline the generation of electromagnetic fields by a non-conformal coupling function that depends on the inflaton and discuss the behavior of the Fourier mode functions describing the electromagnetic vector potential. Further, we shall arrive at the power spectra of the curvature perturbations and the electromagnetic fields in these cases. In Sec. III, we shall consider a non-conformal coupling function that depends on the kinetic energy of the inflaton. We shall show that, for suitable choices of the parameters involved, such a coupling function leads to the desired scale invariant spectrum for the magnetic field in the ultra slow roll scenarios. In Sec. IV, we shall derive the third order action that describes the interaction between the curvature perturbation and the electromagnetic field for the new choice of the non-conformal coupling function. We shall also arrive at formal expressions for the contributions to the three-point cross-correlation (in Fourier space) between the curvature perturbations and the magnetic fields corresponding to the different terms in the action at the cubic order. We shall also introduce the non-Gaussianity parameter associated with the three-point cross-correlation of interest. In Sec. V, we shall calculate the three-point cross-correlation for two specific cases of pure ultra slow roll inflation and suitable values of the relevant parameters. We shall also illustrate the complete structure of the non-Gaussianity parameter and discuss the validity of the consistency condition governing the parameter in the scenarios we consider. Lastly, in Sec. VI, we shall conclude with a summary and outlook. We relegate complementary details of the computation and arguments on certain related issues to the appendices.
@@ II Background, mode functions and power spectra
In this section, we shall discuss the behavior of the background as well as the Fourier mode functions of the curvature perturbation and the electromagnetic field in ultra slow roll inflation. We shall also discuss the power spectra of the curvature perturbations and the electromagnetic fields that arise in a couple of specific situations.
@@ II.1 Evolution of the background
where $H=\dot{a}/a$ is the Hubble parameter and $V_{\phi}={\mathrm{d}}V/{\mathrm{d}}\phi$. As we mentioned above, a brief epoch of ultra slow roll inflation is usually realized in potentials that contain a point of inflection, i.e. a point wherein the first and the second derivatives of the potential, viz. $V_{\phi}$ and $V_{\phi\phi}={\mathrm{d}}^{2}V/{\mathrm{d}}\phi^{2}$, vanish (see, for instance, Refs. [62, 63, 50]). Note that, when $V_{\phi}=0$, the equation of motion for the scalar field reduces to
@@
where $\epsilon_{1\mathrm{i}}$ is the initial value of the first slow roll parameter when the value of the scale factor is $a_{\mathrm{i}}=-1/(H_{{}_{\mathrm{I}}}\,\eta_{\mathrm{i}})$ at the conformal time $\eta_{\mathrm{i}}$. In our discussion below, we shall consider two cases wherein $p=6$ and $p=4$. As we shall see, the first of these leads to a scale invariant spectrum for the curvature perturbations, while the second leads to simple mode functions, which allows us to easily calculate the two and three-point functions of interest [^1] We should clarify a point on nomenclature here. We find that scenarios with constant $p$ are also referred to as constant roll inflation [65, 64]. We shall be focusing on scenarios wherein $p>3$. In such cases, the first slow roll parameter decreases so rapidly that the amplitude of the curvature perturbations grows indefinitely at late times. Hence, we believe it is more apt to refer to these scenarios as ultra slow roll inflation..
@@ II.2 Mode functions describing the curvature perturbation and the scalar power spectrum
There are two points concerning this scalar power spectrum that need to be emphasized. The first point is the fact the spectrum is independent of $k$, i.e. it is scale invariant. The second point is that, in contrast to slow roll inflation, the spectrum does not approach a constant value at late times [i.e. when $k\ll\sqrt{z^{\prime\prime}/z}$, which is equivalent to the super-Hubble limit $k/(a\,H)\simeq(-k\,\eta)\ll 1$]. In fact, the power spectrum grows as $a^{6}$ at late times. This can be attributed to the fact that the mode function $f_{k}$ grows as $a^{3}$ in the super-Hubble limit. Such a growth is a well known feature in the ultra slow roll inflationary scenarios [57, 58].
@@ II.3 Mode functions describing the electromagnetic field and power spectra
We can arrive at the equation of motion governing the vector potential $A_{\mu}$ by varying the above action. In what follows, we shall choose to work in the Coulomb gauge wherein $A_{0}=0$ and $(g^{ij}\,\partial_{i}A_{j})=0$. Let $\bar{A}_{k}$ denote the Fourier mode functions associated with the vector potential $A_{i}$. We find that the mode functions $\bar{A}_{k}$ satisfy the following differential equation (see, for example, Refs. [37, 5]):
@@
where $|0\rangle$ denotes the vacuum state annihilated by the annihilation operator $\hat{b}_{{\bm{k}}}^{\lambda}$, i.e. $\hat{b}_{{\bm{k}}}^{\lambda}|0\rangle=0$ for all ${\bm{k}}$ and $\lambda$. On utilizing the decomposition of the vector potential in Eq. (20), the power spectra $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ and $\mathcal{P}_{{}_{\mathrm{E}}}(k)$ can be expressed in terms of the mode functions $\bar{A}_{k}$ and their scaled forms ${\mathcal{A}}_{k}$ as follows [37, 5]:
@@
where $\eta_{\mathrm{e}}$ is the conformal time at the end of inflation and $\bar{n}$ is a number. Note that the overall coefficient has been chosen so that the coupling function $J$ reduces to unity at the end of inflation (i.e. at $\eta_{\mathrm{e}}$). In situations wherein the scale factor can be expressed in the de Sitter form, the coupling function is given by
@@
For such a choice of the coupling function, the solution to Eq. (18) that satisfies the Bunch-Davies initial conditions in the domain wherein $k\gg\sqrt{J^{\prime\prime}/J}$ [which corresponds to the sub-Hubble limit $(-k\,\eta)\ll 1$ for the $J$ we are working with] can be expressed as
@@
The spectra of the electromagnetic fields can be evaluated at late times when $k\ll\sqrt{J^{\prime\prime}/J}$, i.e. in the super-Hubble limit wherein $(-k\,\eta)\ll 1$. Upon substituting the mode functions given in Eq. (27) in the expressions in Eqs. (24) and considering the super-Hubble limit, the spectra of the magnetic and electric fields $\mathcal{P}_{{}_{\mathrm{B}}}(k)$ and $\mathcal{P}_{{}_{\mathrm{E}}}(k)$ can be obtained to be [37, 5]
@@
$m=\begin{cases}\bar{n},&\text{for $\bar{n}<-\frac{1}{2}$},\\ -\bar{n}-1,&\text{for $\bar{n}>-\frac{1}{2}$},\end{cases}$ (30)
@@
$m=\begin{cases}\bar{n},&\text{for $\bar{n}<\frac{1}{2}$},\\ 1-\bar{n},&\text{for $\bar{n}>\frac{1}{2}$},\end{cases}$ (31)
@@
There are a few points that we need to clarify regarding the results that we have arrived at above. To begin with, note that, we obtain a scale invariant spectrum for the magnetic field when $\bar{n}=2$ and when $\bar{n}=-3$. However, when $\bar{n}=-3$, the spectrum of the electric field behaves as $\mathcal{P}_{{}_{\mathrm{E}}}(k)\propto(-k\,\eta_{\mathrm{e}})^{-2}$, which grows indefinitely at late times. In other words, the energy density in the electric field becomes large leading to significant backreaction on the background (in this regard, see, for instance, Refs. [40, 55]). Therefore, when we require a scale invariant spectrum for the magnetic field, we shall focus on the $\bar{n}=2$ case. Secondly, recall that the original non-conformal coupling function $J$ depends on $\phi$ [cf. Eq. (16)]. Whereas, our choices for $J$ in Eqs. (25) and (26) depend on the scale factor and conformal time. Such a behavior for $J(\eta)$ is achievable in most models of inflation that permit slow roll [39, 38, 76]. However, there exist situations, in particular, when an inflationary potential permits a period of ultra slow roll, wherein a behavior such as $J\propto a^{2}\propto\eta^{-2}$ may not be achievable (for discussions in this context, see Ref. [55]). To circumvent such challenges, either, one has to consider inflation involving more than one field (for a discussion on this point, see Ref. [56]) or, as we shall discuss in the following section, consider more non-trivial forms for the coupling functions involving the kinetic energy of the inflaton.
@@ III Construction of the non-conformal coupling function in pure ultra slow roll inflation
where $J_{0}$ is a constant which we shall choose so that $J$ reduces to unity at the end of inflation. It should be evident that, since the non-conformal coupling function $J(X)$ depends only on time, the electromagnetic field described by the action in Eq. (34) can be treated in the same manner as in the case of the field described by the earlier action in Eq. (16). For instance, if we work in the Coulomb gauge, the Fourier modes $\bar{A}_{k}$ of the vector potential $A_{i}$ will continue to satisfy the same equation of motion, viz. Eq. (17). Also, on quantization, we can elevate the vector potential $A_{i}$ to the operator $\hat{A}_{i}$ as in Eq. (20). Note that, in the ultra slow roll inflationary scenarios of interest, we can write the coupling function in Eq. (35) as
@@
On varying the action in Eq. (34) with respect to the metric tensor, we obtain the stress-energy tensor for the electromagnetic field to be
@@
We find that, in the background described by the FLRW line-element described by Eq. (1), the energy density of the electromagnetic field, viz. $\rho_{{}_{\mathrm{EM}}}=-T^{t}_{t}$, can be expressed as
@@
On using these solutions in Eqs. (24), we obtain the power spectra of the magnetic and electric fields to be
@@ IV.1 Action at the cubic order
As is often done, we shall compute the three-point function using the standard methods of perturbative quantum field theory (for calculations involving scalars and tensors, see, for example, Refs. [96, 97, 98, 99, 100], and for discussions involving magnetic fields, see Refs. [89, 90, 91, 92, 93, 94, 101, 95]). Our first task is to obtain the interaction Hamiltonian at the cubic order that governs the cross-correlation of interest. We shall start with the action in Eq. (34), take into account the scalar perturbations in the spatially flat, FLRW line-element, and arrive at the cubic order action involving the curvature perturbations and the electromagnetic fields. We shall make use of the Arnowitt–Deser–Misner (ADM) formalism to obtain the action [102]. Recall that, in the ADM formalism, the line-element associated with a generic spacetime is described in terms of the lapse function $N$, the shift vector $N^{i}$ and the spatial metric $g_{ij}$ as follows:
@@
Note that the metric elements $g_{00}$ and $g_{0i}$ associated with this line-element are given by
@@
$N_{1}=\frac{{\mathcal{R}}^{\prime}}{a\,H},\;\chi=-\frac{{\mathcal{R}}}{a\,H}+\epsilon_{1}\,\nabla^{-2}{\mathcal{R}}^{\prime}.$ (45)
On substituting the above components of the metric tensor in the action in Eq. (34), we find that the action up to the cubic order involving the curvature perturbation ${\mathcal{R}}$ and the electromagnetic vector potential $A_{i}$ can be expressed as
@@
There are two clarifying remarks that we wish to make regarding the interaction Hamiltonian $H_{\mathrm{int}}$ we have obtained. Firstly, note that, in the above Hamiltonian, if we ignore the terms containing $n$ within the curly brackets, it reduces to the standard form associated with the original action in Eq. (16) involving $J(\phi)$ (in this regard, see, for example, Refs. [90, 91]). Secondly, as with any action at the cubic order, the interaction Hamiltonian proves to be negative of the corresponding cubic order Lagrangian.